# A Study of the Generalized Naming Game and Bayesian Naming Game as Dynamical Systems

### v1

@@ III The Bayesian Naming Game as a Dynamical System
Note that finding a solution to the above system is not possible without making some assumptions about the probabilities $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$, both falling within the interval [0,1] with an unknown functional form. In Ref.Marchetti et al. [2020b], it was proposed to set $p_{A}=p_{B}=p$, thus making the probabilities time-independent. This assumption, along with the initial condition $I_{0}$, played a crucial role in estimating an upper bound of about $20\%$ for the number of agents with two names in their inventories Marchetti et al. [2021]  [^2] The agents having two names in their inventories are often called bilinguals in literature.. However, it cannot be used at all to demonstrate the existence of the fixed points A and B in full generality, as wrongly assumed in Ref.Marchetti et al. [2020b]. In fact, consensus cannot be reached in such a setup due to the absence of symmetry breaking, as previously discussed within the MNG in Section I.
That being said, we cannot infer the exact time-dependent probability pattern of $p_{A},p_{B}$ directly from the simulations of BNG, because it is strongly affected by the system size $N$. However, there is useful information in them, independent of $N$, that can be advantageously exploited. Indeed, it is found that $p_{A},p_{B}$ grow monotonically with time as the success rate observable Marchetti et al. [2020b]. Furthermore, name learning is a few-shot process, where agents successfully generalize from a small number of positive examples Tenenbaum [1999]; Marchetti et al. [2021]. Therefore, the functional form of $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$ cannot be too smooth. With the above considerations in mind, we propose modeling both $p_{A}$ and $p_{B}$ as time-dependent logistic function (or sigmoid function) Murphy [2012]. The logistic function has a S-shaped form that can be modified appropriately thanks to its two characteristic parameters. Denoting it by the symbol $g$, this function reads:
$g\left(t\right)=\frac{1}{1+\exp\left(-\eta\left(t-t_{m}\right)\right)}\,,$ (12)
where $\eta$ is the logistic growth rate (or the steepness of the curve), and $t_{m}$ corresponds to the value of $t$ of function’s midpoint.
Next, we have to tune the parameters of the logistic functions to ensure that the Bayesian probabilities yield the right solution for the problem at hand. In this regard, we need to examine how the Bayesian probabilities behave in multi-agent simulations, assuming that at the beginning $50\%$ of the agents know the word A and the remainder know the other word B Marchetti et al. [2020b]  [^3] From the dynamical system point of view this setup corresponds exactly to have the initial condition $I_{0}$.. In Fig. 7, we plot the cumulative distribution functions (CDF) from the values of $p_{A}$ and $p_{B}$ obtained from a single run of a system with size $N=3,000$ with the above initial condition. Examining these curves derived from simulation data, it is evident that the probabilities $p_{A}$ and $p_{B}$ cannot be assumed to be equal throughout the entire time-evolution of the system, but it may be possible only for a short time interval (see discussion below).
Furthermore, we note that what diversifies the cumulative distribution function (CDF) is the initial bias introduced in the simulations. In order to distinguish the synonyms A and B, different thresholds, $n^{\ast}{ex,A}=5$ and $n^{\ast}{ex,B}=6$, of the number of positive examples stored by the agents, necessary for computing the probabilities $p_{A}$ and $p_{B}$, respectively, are assigned. As a result, this bias will favor the learning of name A. There is also a step-like progression toward unity in the cumulative distribution functions as a direct consequence of few-shot learning Tenenbaum [1999].
@@
Figure 7: Cumulative distribution functions for the values of $p_{A},p_{B}$ obtained from a simulation of the system with size $N=3,000$, starting from the initial condition $I_{0}$ and with Erlang prior. Other simulation parameters are $n^{\ast}_{ex,A}=5$, $n^{\ast}_{ex,B}=6$. Due to this setup, the consensus is reached at A.
That being said, we need to tune the parameters of two logistic functions, here denoted by $g_{A}$ and $g_{B}$, in order to differentiate the time-dependence of probabilities $p_{A}$ and $p_{B}$, respectively. For example, we can set the same value $\eta=2$ for both logistic functions, but at least the parameter $t_{m}$ should have different values to distinguish $g_{A}$ from $g_{B}$. To this end, we set $t_{m,A}=10$ and $t_{m,B}=130$ for $g_{A}$ and $g_{B}$, respectively (see inset of Fig. 8). This parameter choice ensures a step-like pattern for the probabilities, making them non-smooth, while also favoring the learning of name A. With these parameters in place, we expect that the time-evolution of the dynamical system reproduces the main features observed in simulations of multi-agent system.
To confirm the above conjecture we proceed to numerically integrate Eqs. 10, 11, replacing $p_{A}$ with $g_{A}$ and $p_{B}$ with $g_{B}$. The numerical result shows that the solution $\left(x,y\right)$ asymptotically converge to $A=\left(1,0\right)$ (see Fig. 8444We attempted to numerically integrate the differential system using both the Runge-Kutta method and the LSODA algorithm. We found that both methods produced identical solution.. Furthermore, by swapping the values of parameters $t_{m,A}$ and $t_{m,B}$, it is also straightforward to show that there exist the consensus at $B=\left(0,1\right)$. Overall, there exist a spontaneous emergence of consensus despite the dynamics follows a mean-field deterministic prescription.
Remarkably, the numerical integration also yields a solution $\left(x,y\right)$ that closely resembles that obtained from a simulation with relatively large number of agents, i.e., $N\gtrsim 200$ (see Ref. Marchetti et al. [2020b]).
@@
We conclude noting that the solution $\left(x,y\right)$ splits in its two components $x$ and $y$ after a transient time during which $x=y$, see Fig. 8. Such an initial stage of dynamics, where $x=y$, can be thought as a dynamical system whose differential equations are: $\dot{x}=-p_{B}xy$, $\dot{y}=-p_{A}xy$. This fact clearly explains that the dynamics of BNG dynamical system is primarily dictated by the Bayesian probabilities $p_{A}$ and $p_{B}$, at least in the initial stage of the dynamics. Once again, it is their interplay and time-dependence that lead to the subsequent symmetry-breaking process, essential for the emergence of consensus.
Figure 8: Solution of Eqs. 10, 11 assuming $g_{A}$ and $g_{B}$ with parameters $\eta=2$, $t_{m,A}=10$ and $t_{m,B}=130$, The differential equations at hand were numerically integrated by means of LSODA algorithm assuming $I_{0}=\left(0.5,0.5\right)$.
@@ Acknowledgments
The author would like to express gratitude to Steven Strogatz for his helpful insights on the bifurcation and to Konstantinos “Kostas” Zygalakis for the important remarks about the effect of noise around the critical points and bifurcations and Andrea Baronchelli for reading the manuscript. Some results presented here were obtained during the Spring of 2021 while the author was at the National Institute of Chemical Physics and Biophysics, Tallinn, Estonia. Therefore, the author extends thanks to Martti Raidal for the financial support via the European Regional Development Fund CoE program TK133 “The Dark Side of the Universe”, and to Els Heinsalu and Marco Patriarca for proposing the naming game as a research topic.

### v3

@@ IV The Bayesian Naming Game as a Dynamical System
Note that finding a solution to the above system is not possible without making some reasonable assumptions about the probabilities $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$, both falling within the interval [0,1] with an unknown functional form. In Ref.Marchetti et al. [2020b], it was proposed to set $p_{A}=p_{B}=p$, thus making the probabilities identical and time-independent at the same time. While this assumption, along with the initial condition $I_{0}$, played a crucial role in estimating an upper bound of about $20\%$ for the number of agents with two names in their inventories Marchetti et al. [2021]  [^5] The agents having two names in their inventories are often called bilinguals in the literature., it cannot be used to demonstrate the existence of fixed points A and B in full generality, as wrongly assumed in Ref.  Marchetti et al. [2020b]. In fact, consensus cannot be reached in such an initial setup due to the absence of spontaneous symmetry breaking Goldstone [1961]; Michel [1980]  [^6] Here the term “spontaneous” indicates that the symmetry breaking occurs without the intervention of an external agent, similar to what happens with the Higgs mechanism in particle physics., similar to what happens to the dynamical system associate with MNG (see Section I).
Next, we shall demonstrate that the assumption $p_{A}=p_{B}=p$ cannot be valid throughout the entire time-evolution of the dynamical system associated with BNG. To illustrate this fact we computed cumulative distribution function (CDF) of the values taken by $p_{A}$ and $p_{B}$ during a single multi-agent simulation of a system with size $N=3,000$ (see Section II for details).
@@
Figure 6: The cumulative distribution function of $p_{A}$ (solid line) and $p_{B}$ (dashed line) as a function of the Bayesian probabilities $p_{A}$ and $p_{B}$, respectively, obtained from a simulation of a system with size $N=3,000$, starting from the initial condition $I_{0}$, and assuming the Erlang prior. The name learning thresholds for A and B, were $n^{\ast}_{ex,A}=5$ and $n^{\ast}_{ex,B}=6$, respectively. In such a case the consensus is reached at A.
In Fig. 6, the curves corresponding to CDF of $p_{A}$ (solid line) and $p_{B}$ (dashed line) are shown as functions of the Bayesian probabilities. From their different trends, it is clear that the assumption $p_{A}=p_{B}=p$ is untenable and may only be valid for a short time-interval. Furthermore, there is also a step-like progression toward unity in the two cumulative distribution functions as a direct consequence of the few-shot learning processes Tenenbaum [1999a]. The latter feature of this model demonstrates that it is implausible to consider such probabilities as smooth functions either of the number of positive examples or of time. Regarding their time-dependence, we cannot infer its exact functional form directly from the multi-agent simulations, due to their strong dependence on the system size $N$. However, Bayesian probabilities exhibit an evident monotonically increasing trend with time within a mean-field approximation Marchetti et al. [2020b].
With the above considerations in mind, we propose modeling both $p_{A}\left(t\right)$ and $p_{B}\left(t\right)$ as a time-dependent logistic function (or sigmoid function) Murphy [2012]  [^7] The logistic function is a common choice of activation function for the neural networks and deep learning.. The logistic function has a S-shaped form that can be modified appropriately thanks to its two characteristic parameters: the logistic growth rate (or the steepness of the curve) $\eta$, and the quantity $t_{m}$ that corresponds to the value of $t$ of function’s midpoint. Denoting it by the symbol $g$, the logistic function reads:
$g\left(t\right)=\frac{1}{1+\exp\left(-\eta\left(t-t_{m}\right)\right)}\,.$ (12)
In the following, we shall give an example of how our probabilistic modeling yields reasonable results in accordance with the numerical simulations. Without the loss of generality, we chose the following three values for parameters $\eta$ and $t_{m}$: $\eta=2$, $t_{m,A}=10$ and $t_{m,B}=130$. Inserting $\eta=2,t_{m,A}=10$ and $\eta=2,t_{m,B}=130$ into Eq. 12 we get two different logistic functions: $g_{A}$ and $g_{B}$, respectively. While such a value of $\eta$ guarantees that both $g_{A}$ and $g_{B}$ are step-like non-smooth functions of time $t$, the different values of $t_{m}$ i.e. $t_{m,A}$, $t_{m,B}$, differentiate $g_{A}$ from $g_{B}$. Furthermore, our choice of parameters ensures that the learning of name A will be favored.
Figure 7: Numerical solution of Eqs. 10, 11 replacing the probabilities $p_{A}$ and $p_{B}$ with the logistic functions $g_{A}$ and $g_{B}$, respectively. The logistic function’s parameters used are $\eta=2$ for both $g_{A}$ and $g_{B}$, $t_{m,A}=10$ and $t_{m,B}=130$ for $g_{A}$ and $g_{B}$, respectively. The initial condition is $I_{0}=\left(0.5,0.5\right)$.
Next, we proceed to numerically integrate Eqs. 10, 11, replacing $p_{A}$ with $g_{A}$ and $p_{B}$ with $g_{B}$, using the algorithm LSODA Petzold [1983], and assuming the initial condition $I_{0}=\left(0.5,0.5\right)$. The numerical result shows that the solution $\left(x,y\right)$ asymptotically converges to $A=\left(1,0\right)$ (see Fig. 7). Accordingly, there exists now a spontaneous emergence of consensus despite the dynamics following a mean-field deterministic prescription. Furthermore, by swapping the values of parameters $t_{m,A}$ and $t_{m,B}$, it is also straightforward to show that the solution converges to the other possible consensus state, that is, $B=\left(0,1\right)$. These outcomes are attributed to the spontaneous symmetry breaking induced by the time-dependent probabilities, disrupting the invariance under the swap of variables $x$ and $y$ of Eqs. 10 and  11. Remarkably, the numerical integration yields a solution $\left(x,y\right)$ that closely resembles that obtained from the numerical simulations with relatively large number of agents, i.e., $N\gtrsim 200$ (see Ref. Marchetti et al. [2020b]).
@@
We conclude by noting that the numerical solution $\left(x,y\right)$ splits into its two components $x$ and $y$ after a transient time during which $x=y$ (see Fig. 7). Such an initial stage of dynamics, can be thought as dictated by a dynamical system, whose differential equations read: $\dot{x}=-p_{B}xy$, $\dot{y}=-p_{A}xy$. This observation clearly explains that in the initial stage of the dynamics, the time-evolution of the system is now primarily dictated by the Bayesian probabilities $p_{A}$ and $p_{B}$. Once again, it is their interplay and time-dependence that lead to the subsequent spontaneous symmetry-breaking process, essential for the emergence of consensus.
@@ Acknowledgments
The author would like to express gratitude to Steven Strogatz for his helpful insights on bifurcation, to Konstantinos ”Kostas” Zygalakis for the important remarks about the effect of noise around critical points and bifurcations, and to Andrea Baronchelli for reading the manuscript. The author extends thanks to Martti Raidal for the financial support via the European Regional Development Fund CoE program TK133 “The Dark Side of the Universe”, and to Els Heinsalu and Marco Patriarca for proposing the naming game as a research topic.
Appendix A Sample Variance of the Random Variable $X_{t}$
The theory of stochastic processes provide the following facts about the variance $Var[X_{t}]$ of Brownian motion and OU process. For the former it is expected that the variance grows linearly with time, i.e., $Var[X_{t}]\propto\sigma_{1}t$ ($\sigma_{1}$ is a suitable positive parameter) while for the latter is given by the following equation Oksendal [2010]:
$Var[X_{t}]=\frac{\sigma_{2}^{2}}{2\kappa}\left(1-e^{-2\kappa t}\right)\,,$ (13)
where $\sigma_{2}$ and $\kappa$ are two suitable parameters. In Fig. 8 we plot the sample variances of the random variable $X_{t}$ as function of time $t$, computed from $1,200$ trajectories of the $\beta$-NG model, assuming $\beta=0.2$ (black line) and $\beta=\beta_{c}$ (blue line), respectively. It is evident that while the curve corresponding to $\beta=\beta_{c}$ exhibits the expected linear trend of a Brownian motion, the other one corresponding to $\beta=0.2$ correctly behaves according to Eq. 13. To quantitatively support our observations, we proceed to fit the data to the above formulas. The curve fitting procedure by non-linear least squares yields the following optimal values for the parameters: $\sigma_{1}=0.00016\pm 1.07\times 10^{-8}$ with $R^{2}\approx 0.99$ and $\sigma_{2}=0.013\pm 7.7\times 10^{-6}$, $\kappa=0.132\pm 1.5\times 10^{-4}$ with $R^{2}\approx 0.93$ where $R^{2}$ denotes the coefficient of determination. These values of $R^{2}$ indicate a good fit to the data, thereby further supporting our considerations regarding the type of stochastic processes for the problem at hand.
Figure 8: Sample variance $Var[X_{t}]$ of the random variable $X_{t}$ ($X_{0}=0.5$) as function of time $t$. The variance is computed from $1,200$ trajectories of the $\beta$-NG model, assuming $\beta=0.2$ (black line) and $\beta=\beta_{c}$ (blue line), respectively. Note that for a better visualization we multiply both sample variances by a factor $10$.